# Debiasing Welch’s Method for Spectral Density Estimation

### 21 Dec 2023 v1

@@ 6 Unevenly spaced bases for signal compression
To be analogous to Welch’s estimator, in Section 3 we define the debiased Welch estimator with even basis spacing over frequency. As noted therein, we do not require this, and can alternatively define the bases by an irregular partition over frequency. For spectra whose value varies slowly over certain neighbourhoods of frequency, we argue that a wider basis spacing would lead to little reduction of accuracy in estimating the true process. Allowing for wider, and irregularly spaced bases serves two purposes: first we achieve a degree of signal compression, as compared to the Welch estimator; and second, we reduce the variance of the estimator as there are more data informing the coefficient of each basis. The code provided alongside this manuscript accepts uneven bases by providing the $\omega^{\mathrm{c}}_{k}$ and $\delta_{k}$, as defined in (13), as user defined inputs for all $k$.
To demonstrate this concept, we calculate Welch and debiased Welch estimates of a random sample from a Matérn process parameterized as
$\gamma(\tau)=\frac{2\sigma^{2}}{\Gamma(\alpha-1/2)2^{\alpha-1/2}}|\tau|^{\alpha-1/2}\kappa_{\alpha-1/2}(|\tau|),\hskip 14.22636ptf(\omega)=\frac{\sigma^{2}}{(\omega^{2}+\lambda^{2})^{\alpha}},$
with $\sigma=1$, $\lambda=0.1$ and $\alpha=11/6$, where $\Gamma(\cdot)$ denotes the gamma function and $\kappa_{\alpha-1/2}(\cdot)$ the order $\alpha-1/2$ modified Bessel function of the second kind. The true power spectral density and the expected value of Welch’s estimator for the Matérn process are shown in the plots in Figure 5 by the dashed and solid grey lines, respectively. The Matérn power spectral density is asymptotically, in $\omega$, log-linear and so we space the bases of the debiased Welch estimator evenly in log-frequency. The location of the bin-centres, $\omega_{k}^{\mathrm{c}}$, are shown by the rug-plot on the frequency axes in Figure 5. Asymptotically log-linear processes are common in many sciences, for example, background ocean spectra and turbulent dissipation in oceanography (Garrett and Munk, 1972; Zulberti et al., 2022), stellar rotation periods and red-shift in astronomy (Foreman-Mackey et al., 2017) and long-memory processes (Palma, 2007), and so this example provides a useful case-study to many applied fields. We sample a length $32,768$ $(2^{15})$ time series with $\Delta=1$ and calculate the Welch estimate, with parameters $M=32$ $(2^{5})$, $L=1024$ $(2^{10})$ and no overlap, shown in the left plots of Figure 5 in black. The debiased Welch estimate, with log-spaced bases, is calculated and shown in the right plots of Figure 5 in black. The bottom plots of Figure 5 are identical to the top, except with frequency shown on a log scale. Not only do we reduce empirical bias in the estimate, we also compress the signal by a factor of $\sim 50$ and reduce the standard deviation of the estimate from the true spectral density.
@@
Figure 5: Welch estimates of a Matérn power spectrum with uneven basis spacing. The top plots show frequency on a linear scale and the bottom plots on a log scale. The standard Welch estimate is shown left and debiased Welch estimate, with uneven basis spacing, is shown right with black solid lines. Locations of the Fourier frequencies and basis centres for the Welch and debiased Welch estimates are shown by the rug plots on the frequency axes. In all plots, the true power spectral distribution and expected Welch estimate are shown by the dashed and solid grey lines, respectively.

### 11 Apr 2024 v2

@@ 6 Unevenly spaced bases for signal compression
Figure 6: Welch estimates of a Matérn power spectrum with uneven basis spacing. The top plots show frequency on a linear scale and the bottom plots on a log scale. The standard Welch estimate is shown left and debiased Welch estimate, with uneven basis spacing, is shown right with black solid lines. Locations of the Fourier frequencies and basis centres are shown by the rug plots on the frequency axes. In all plots, the true power spectral distribution and expected Welch estimate are shown by the dashed and solid grey lines, respectively.
To be analogous to Welch’s estimator, in Section 3 we define the debiased Welch estimator with even basis spacing over frequency. As noted therein, we do not require this, and can alternatively define the bases by an irregular partition over frequency. For spectra whose value varies slowly over certain neighbourhoods of frequency, we argue that a wider basis spacing would lead to little reduction of accuracy in estimating the true process. Allowing for wider, and irregularly spaced bases serves two purposes: first we achieve a degree of signal compression, as compared to the Welch estimator; and second, we reduce the variance of the estimator as there are more data informing the coefficient of each basis. The code provided alongside this manuscript accepts uneven bases by providing the $\omega^{\mathrm{c}}_{k}$ and $\delta_{k}$, as defined in (13), as user defined inputs for all $k$. To demonstrate this concept, we return to the Matérn model simulated in Section 4.2. We sample a length $32,768$ $(2^{15})$ time series and calculate the Welch estimate, with parameters $M=32$ $(2^{5})$, $L=1024$ $(2^{10})$ and no overlap, shown in the left plots of Figure 6 in black. The true power spectral density and the expected value of Welch’s estimator for the Matérn process are shown in the plots in Figure 6 by the dashed and solid grey lines, respectively. The Matérn power spectral density is asymptotically, in $\omega$, log-linear and so we space the bases of the debiased Welch estimator evenly in log-frequency. The location of the bin-centres, $\omega_{k}^{\mathrm{c}}$, are shown by the rug-plot on the frequency axes in Figure 6. The debiased Welch estimate, with log-spaced bases, is calculated and shown in the right plots of Figure 6 in black. The bottom plots of Figure 6 are identical to the top, except with frequency shown on a log scale. Not only do we reduce empirical bias in the estimate, we also compress the signal by a factor of $\sim 50$ and reduce the standard deviation of the estimate from the true spectral density.